Session: Nov 2003 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Urumqi (yangwj at Thu Oct 30 15:49:24 2003) Minor failures : RECE TSYS LATE
The cooling receiver is fail,so we changed to use old non-cooling useful recording start at UT 12:11 .T-cal values in log file are wrong, real T-cal value is 174.5K.

Medicina (A.Maccaferri at Thu Oct 30 18:47:36 2003) Minor failures : FS
Incorrect setting of IFD at start, corrected after few seconds from start, so first tsys for vc02 and vc04 are wrong.

Torun (K.Borkowski at Fri Oct 31 06:59:09 2003) Failure : RECO
Our P&G recorder is still unusable; the struggle continues.

Hartebeesthoek (Jonathan Quick at Fri Oct 31 07:41:25 2003) Success :

Onsala (Michael Lindqvist at Fri Oct 31 07:58:40 2003) Success :
No known problems.

Effelsberg (A. Kraus at Fri Oct 31 09:11:22 2003) Success :

Jodrell_Bank (Alastair Gunn at Mon Nov 3 10:37:46 2003) Success :
No known problems.

Noto (P. Cassaro at Mon Nov 3 13:22:58 2003) Success :

Westerbork (Tony Foley at Tue Nov 4 14:47:15 2003) Success : RECO
This was recorded onto mark5 as the capstan motor of the tape drive was being repaired using cannibalized parts from an old Honeywell drive